provta that (sin239-sin221)(sin257-sin23)=3/16

provta that
(sin239-sin221)(sin257-sin23)=3/16

Grade:

1 Answers

Sumit Majumdar IIT Delhi
9 years ago
Solution:
We have:
sin(239)-sin(221)=2cos\left (\frac{239+221}{2} \right )sin\left (\frac{239-221}{2} \right )=2cos(230)sin(9)
similarly,
sin(257)-sin(23)=2cos\left (\frac{257+23}{2} \right )sin\left (\frac{257-23}{2} \right )=2cos(140)sin(117)
Simplifying this should give the required answer.
Regards
Thanks & Regards
Sumit Majumdar,
askIITians Faculty
Ph.D,IIT Delhi